Chapter 8 In class assignment- submitted in Canvas

Use these factors to answer questions

n

i

PV$1

FV$1

PVA

FVA

3

5%

0.86384

1.15763

2.72325

3.1525

5

6%

0.74726

1.33823

4.21236

5.6371

60

0.50%

0.74137

1.34885

51.72556

69.7700

  1. You want to have $30,000 at the end of 5 years for a down payment on a house.  How much should you save each month to achieve your goal if you can earn 6%? $______ (round to nearest dollar)

  1. You want to buy a $30,000 car today and will be making monthly payments for the next 5 years.  What is your car payment if your borrowing rate is 6%? $______ (round to nearest dollar)

  1. You won the local lottery for $60,000.  Your choice is to take the $60,000 today or if you wait you can get $67,000 bonus in 3 years (guaranteed).  Your current investment rate is 5%.  What is the best economic decision, take the amount today or wait 3 years?
  1. You plan to invest $2,000 every year for the next 5 years.  What will be the value of your investment at the end of the 5 years if you can earn 6% annual interest?  $_______ (round to nearest dollar)

  1. Your parents have decided to give you $20,000 for a down payment on a house in 5 years.  Assuming your parents are earning 6% annually, how much do they need to invest today so they have the money to give you in 5 years?  $______ (round to nearest dollar)
Answer #1

1.as we need to find the present value for 30000$ at the rate of 6% for five years

we will use PV factor

=30000*0.74726

=22418$

2.we will use Future value to find out the value of 30000$ vcar after 5 years at 6%

30000*1.33823

=40147$

3.future value of 60000$ n 3 r=5%

60000*1.15763

=69458$

after 3 years you are getting67000$ which is lesser

so the economic deicion is to accept money at present

4.2000$ is invested every year so will useFVA

=2000*5.6371

=11274$

5. present value of 20000$

=20000*0.74726

=14945$